OpenStudy (anonymous):

Please help if you are fluent with the terms cos, etc. When you get your answer could you please explain to me how you got it? It would mean so much. Thank you, in advance.

OpenStudy (anonymous):

Here's the equation. \[4 \cos \theta = 3 \cos \theta -\sqrt{3/2} for 0 \le \theta <2\]

OpenStudy (anonymous):

And here are my possible answers. A. \[5\pi/6, 7\pi/6\] B. \[\pi/6, 11\pi/6\] C. \[\pi/3, 5\pi/3\] D. none of these.

OpenStudy (kc_kennylau):

Take 3cos(theta) from both sides :)

OpenStudy (kc_kennylau):

(take=subtract)

OpenStudy (anonymous):

Alright, so factor them out. :)

OpenStudy (kc_kennylau):

not factor :p

OpenStudy (anonymous):

in what quadrants is cosine a negative value?

OpenStudy (kc_kennylau):

\[\begin{array}{rcl} 4\cos\theta&=&3\cos\theta-\frac{\sqrt3}2\\ 4\cos\theta-3\cos\theta&=&-\frac{\sqrt3}2\\ \cos\theta&=&-\frac{\sqrt3}2 \end{array}\]

OpenStudy (anonymous):

@kc_kennylau yes, I understand that.

OpenStudy (anonymous):

Thank you :) what happens next?

OpenStudy (anonymous):

In which quadrants is cosine negative?